OpenStudy (anonymous):

Solve the equation. 5/13 t = -9

OpenStudy (kira_yamato):

t = -117/5

OpenStudy (anonymous):

(5 t)/13 = -9 .......t=-117/5

OpenStudy (anonymous):

what r the options??

OpenStudy (anonymous):

r u trying to solve it?

OpenStudy (anonymous):

(5 t)/13 = -9